College students take different course loads. Many colleges determine the number of 'Full-time equivalent students, F, by using the formula F=n/15, where ni is the total number of credits in which a student enroills for a given semester.

If a student on a campus registers for a total of 41670 credits, how many full-time equivalent students does the campus have?
The campus has _________ equivalent students.
Solve the formula for n.
n=

Answer by nyc_function(2741)

I assume this question has two parts.
We given a formula and what the letters represent in the formula.
For part (a), we are solving for F.
F = n/15
F = 41670/15
F = 2778 equivalent students
======================================
For part (b), we are solving for n. This means we need to isolate n on one side of the equation.
F = n/15
Multiply both sides of the equation by 15/1 to clear the fraction on the right side.
F * 15/1 = n/15 * 15/1
15F = n
Did you follow?